OpenStudy (anonymous):

Georgia built a model of the Empire State Building that is proportional to the actual building. The actual building is 187 feet wide and 1,454 feet tall. Her model is 10 feet tall. Approximately how wide is Georgia's model?

OpenStudy (anonymous):

h1/wi=h2/w2

OpenStudy (anonymous):

@theyatin ?

OpenStudy (anonymous):

height of tower/width of tower=height of model/width of the model. that is 1454/187=10/w solve it for w

OpenStudy (anonymous):

@theyatin you know the answer?? lol.